OpenStudy (idani):

Find a quadratic model for the set of values: (-2, -20), (0, -4), (4, -20).

OpenStudy (danjs):

The standard parabola is - y = ax^2 + bx + c you can use the points given with that to figure out what a,b,c are equal to..

OpenStudy (danjs):

(0, -4) -4 = a*0^2 + b*0 + c (4 , -20) -20 = a*4^2 + b*4 + c (-2 , -20) -20 = a*(-2)^2 + b*(-2) + c Those simplify.. -4 = c -20 = 16*a + 4*b + c -20 = 4*a - 2*b + c